OpenStudy (anonymous):

If 2x + 3y = 17 and 3x + 4y = 24, then 5x +7y = 7 41 3 -7 -41

OpenStudy (yanasidlinskiy):

I'll use the first one 2x+3y=17 => 2x=17-3y => x=17-3y/2 now put this new equation into the second one 3x+4y=24 => 3((17-3y)/2)+4y=24 now to work it out. (51-9y)/2+4y=24 so (51/2)-(9y/2) +4y=24 then (-9y/2)+4y=24-(51/2) -4.5y+4y=24-25.5 (-1/2)y=-3/2 y=(-3/2)/(-1/2) y=3 when y=3 find x by putting 3 were y is 3x+4y=24 3x+4(3)=24 3x+12=24 3x+24-12 => 3x=12 therefore x=4 now that y=3 and x=4